How to solve 7x2=21x ? (seven x squared equals twenty-one x). The answer key says the answer is 0,3 but I do not how how they solved it. My educated guess is that I need to make them both equal to zero but am not sure. Thank you for your help and look forward to hearing from you.
Answer by stanbon(75887)

7x2=21x ? (seven x squared equals twenty-one x).
The answer key says the answer is 0,3 but I do not how how they solved it.
================
7x^2-21x = 0
Factor:
7x(x-3) = 0
7x = 0 or x-3 = 0
---
x = 0 or x = 3
========================
Cheers,
Stan H.
===================